OpenStudy (anonymous):

PLEASE HELP! Which equation would best help solve this problem? Five added to 4 times a number is equal to 9 less than two times the number. 4x – 5 = 2x + 9 5x + 4 = 9x – 2 4x + 5 = 2x – 9 9x + 2 = 5x + 4

OpenStudy (anonymous):

c

OpenStudy (anonymous):

How do you know?

OpenStudy (anonymous):

4 times a number is 4x and add 5 is 4x+5 and then 2 times a number is 2x and 2x-9 is 9 less than 2x.

OpenStudy (anonymous):

Okay, what about this one. Hold on

OpenStudy (anonymous):

Find the unknown number. Six added to 3 times a number is equal to 4 less than 5 times the number. 5 –5 –10 10

OpenStudy (anonymous):

Type it out..Make sure you know how to do it. 3 times a number can be written as 3x. + 6 for six added. You can take it from there.

OpenStudy (anonymous):

3 times a number is represented how? (look at the last answer I gave you) Then add 6 to that. "Is equal to" means to set all that above = to something 5 times a number is what? Then use that like the last example to get "5 times a number" - 4 (this would be 4 less, correct)?

OpenStudy (anonymous):

Oh wait, hold on, LOL.

OpenStudy (anonymous):

ya i think i got it, thanks guys. It was easier than I thought... hahaha.

OpenStudy (anonymous):

:)

OpenStudy (cookielate):

"Five" . . . that one is pretty easy. "4 times a number" . . . 4Q "Five added to 4 times a number" . . . 5 + 4Q "2 times the number" . . . 2Q "9 less than 2 times the number" . . . 2Q - 9 The first quantity is equal to the second quantity . . . 5 + 4Q = 2Q - 9 That's choice D. What if you had to solve it, and find the mysterious unknown number. Could you handle it from here ? 5 + 4Q = 2Q - 9 Subtract 5 from each side: 4Q = 2Q - 14 Subtract 2Q from each side: 2Q = -14 Divide each side by 2 : Q = -7 -7 is the only number that will make that original long messy statement true.